Grade 11General Physics

The acceleration of a particle is increasing linearly with time t as bt. The particle starts from origin with an initial velocity Vo. The distance travelled by particle in time t will be

Aman Singh
7 Years agoGrade 11

1 Answer

Venkat
7 Years ago
a(t) = bt
v(t) = bt2/2+c
Given, c = V0
v(t) = bt2/2+V0
d(t) =  bt3/6+V0t+C
starts from origin means C= 0. Hence,
d(t) = bt3/6+V0t